40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 11 18:20:59 2018, gautam, Update, Optical Levers, BS/PRM HeNe is dead BSOL_dead.png
    Reply  Sat Dec 15 20:04:03 2018, gautam, Update, IOO, TT1/TT2 stepping 
       Reply  Wed Dec 19 15:15:56 2018, gautam, Update, IOO, TT1/TT2 stepping 
Message ID: 14362     Entry time: Sat Dec 15 20:04:03 2018     In reply to: 14345     Reply to this: 14368
Author: gautam 
Type: Update 
Category: IOO 
Subject: TT1/TT2 stepping 

I'm running a script that moves TT1 and TT2 randomly in some restricted P/Y space to try and find an alignment that gets some light onto the TRY PD. Test started at gpstime 1228967990, should be done in a few hours. The IMC has to remain locked for the duration of this test. I will close the PSL shutter once the test is done. Not sure if the light level transmitted through the ITM, which I estimate to be ~30uW, will be enough to show up on the TRY PD, but worth a shot I figure.

Test was completed and PSL shutter was closed at 1228977122.

ELOG V3.1.3-